OpenStudy (madgirlwithabluebox):

write a point slope form of the line through point J(4,1) with slope -4

OpenStudy (anonymous):

Y=mx+b B is what has to be solved for, you know x y and m, plug those in, solve for b Then write the equation and replace m and b with the values you have, but leave x and y in the equation ^_^

OpenStudy (triciaal):

point-slope form for the equation of a line y = mx + b using the point (4, 1) x = 4 and y = 1 and given the slope m = -4 with this solve for b 1 = -4(4) + b b = 1 + 16 = 17 so for any point on this line (x, y) y = -4(x) + 17